### v1

Abstract
Click-fraud is commonly seen as a key vulnerability of pro-rata revenue sharing on music streaming platforms, whereas user-centric is largely immune. This paper develops a tractable non-cooperative model in which artists can purchase fraud activity that generates undetectable fake streams up to a technological limit. We show that pro-rata can be fraud-robust: when fraud technology is weak, honesty is a strict dominant strategy, and an efficient fraud-free equilibrium obtains. When fraud technology is strong, a unique fraud equilibrium arises, yet aggregate fake streams remain bounded. Although fraud is inefficient, the resulting redistribution may improve fairness in some cases. To mitigate fraud without abandoning pro-rata, we introduce a parametric weighted rule that interpolates between pro-rata and user-centric, and characterize parameter ranges that restore a fraud-free equilibrium under technology constraint. We also discuss implications of Spotify’s modernized royalty system for fraud incentives.
Keywords: music streaming platform, pro-rata, user-centric, revenue sharing, strategic manipulation, click-fraud.
JEL Classification: C72, K42, L82
1 Introduction
Music streaming platforms (e.g., Spotify, Apple Music, and Deezer) charge users a subscription fee and share revenue with artists who supply music content, according to users’ streams and the platform’s revenue-sharing rule. Two rules commonly used in practice are pro-rata and user-centric. Under pro-rata, an artist receives a share equal to their streams divided by total streams; under user-centric, each user’s payment is allocated proportionally to the artists that user streams.
In the debate between these two rules, a major concern for pro-rata is click-fraud (Dimont, 2018). The reason is straightforward: if an artist can generate an unlimited number of artificial streams under pro-rata, they could capture almost all royalties. Examples include Vulfpeck’s Sleepify and Michael Smith’s bot accounts, as discussed in the literature (Dimont, 2018; Ghosh et al., 2025). Vulfpeck created an album consisting of silent tracks and asked fans to stream it while asleep. Michael Smith posted AI-generated songs and used thousands of bot accounts to boost streams. In both cases, fraudulent activity increased royalty payments.
Strategic manipulation by artists, i.e., click-fraud, has been studied from an axiomatic perspective. Bergantiños and Moreno-Ternero (2025) propose an axiom called click-fraud proofness, motivated by Vulfpeck’s Sleepify. It requires that an artist cannot gain more than the subscription fee when a user reallocates streams. Ghosh et al. (2025) strengthen this requirement with single-user bribery-proofness, which bounds the bonus to any subset of artists by the subscription fee paid by an additional user. The single-user fraud-proofness axiom in Ghosh et al. (2025) captures the logic behind Michael Smith’s bot accounts: an artist should not profit from faking a new user by more than the subscription cost. User-centric satisfies these axioms, whereas pro-rata does not.
However, this does not mean that pro-rata’s vulnerability to fraud is insurmountable. As noted in Ghosh et al. (2025), these axioms impose no upper bound on streams generated by manipulated users. In practice, even without advanced fraud detection, platforms can detect and remove large volumes of suspicious streams. To capture this idea, we model a fraud technology that produces a fixed number of streams, $\lambda_{0}$, per unit of fraud activity paid by artists. The parameter $\lambda_{0}$ represents the maximum number of streams from each fake user that the platform cannot detect. Interpreted in relative terms, improvements in detection reduce the scope for subtle manipulation, while evolving fraud methods may make manipulation harder to detect.
Previewing the main results, we show that honesty is a strict dominant strategy for all artists when fraud technology is not too strong, with the threshold depending on artists’ real streamshares. Even when fraud technology is strong and cannot be fully controlled, total fake streams remain bounded. Under pro-rata, an increase in total streams dilutes the value of each stream, reducing the marginal benefit from additional fake streams until it falls below the fraud cost. Moreover, fraud may lead to a fairer distribution: it benefits cheaters while reducing the income of high-streamshare honest artists by diluting stream value, potentially increasing the minimum payoff.
We also discuss alternative approaches. First, we extend the analysis to a class of weighted rules that are convex combinations of pro-rata and user-centric. These rules differ in their ability to deter fraud and allow the platform to balance simplicity and robustness, maintaining efficiency under constraints on fraud technology. Second, we discuss Spotify’s recent modernization of its royalty system, focusing on a qualification policy that introduces an eligibility threshold for revenue sharing. We show that, when the qualification threshold is not too high, the policy either fails to deter fraud or harms all artists.
This paper is the first to study strategic manipulation by artists on music streaming platforms using a non-cooperative game. By analyzing pro-rata’s vulnerability to fraud, the paper contributes to the ongoing debate between pro-rata and user-centric. Beyond conceptual discussions and axiomatic analyses (Dimont, 2018; Bergantiños and Moreno-Ternero, 2025; Ghosh et al., 2025), Alaei et al. (2022) and Lei (2023) are the main theoretical contributions in this area. Alaei et al. (2022) study a model with participation decisions by both users and artists. Lei (2023) shows that pro-rata can be more efficient and fairer than user-centric from the users’ perspective when music quality is endogenously chosen. In contrast to these long-run approaches that emphasize participation and quality, our paper focuses on artists’ short-run strategic manipulation, thereby enriching this literature.
Contest technology is applied in analysis. Fraud activity links the pro-rata sharing rule to standard contest models with headstarts. The contest success function in our model is axiomatized by Yu (2025) through “No advantageous reallocation,” extending Amegashie (2006) from the symmetric case to a contest framework with equal starts. Franke et al. (2018) study lottery contests with linear contest technology in a setting that is more general than ours, focusing on the optimal design of headstarts to maximize effort. In our model, effort is purely socially wasteful, and the focus is on equilibrium conditions, bounds on aggregate fraud, and fairness comparison.
The remainder of the paper is organized as follows. Section 2 presents the model. Section 3 analyzes the fraud-free and fraud equilibria and compares fairness. Section 4 discusses alternative solutions. Section 5 concludes.
2 Model
Consider a two-sided music platform that hosts both users and artists. Each user pays a subscription fee normalized to 1, and artists share the $\beta\in(0,1)$ fraction of total revenue. Denote by $N$ the set of $n$ artists and by $M$ the set of $m$ users. For each user $j\in M$, let the total number of streams be $\lambda_{j}>0$, and let $\pi_{ij}\in[0,1]$ denote the share of user $j$’s streams allocated to artist $i\in N$, with $\sum_{i\in N}\pi_{ij}=1$. Thus, the streams from user $j$ on artist $i$ are $\lambda_{j}\pi_{ij}$. The model is considered over a short subscription term (e.g., monthly). Over the short term, artist participation is fixed. Combining full participation of artists with constant user preferences, the set of users maintaining their subscriptions is also unchanged (Alaei et al., 2022). Thus $\lambda_{j}$ and $\pi_{ij}$ are constant for all $i\in N$ and $j\in M$.
Revenue-sharing rules allocate the revenue $m\beta$ to artists based on users’ streams. Under pro-rata, revenue is shared among artists according to their streamshares. The revenue paid to artist $i$ is
$R_{i}^{pr}=\frac{\sum_{j\in M}\lambda_{j}\pi_{ij}}{\sum_{j\in M}\lambda_{j}}m\beta.$
In contrast, user-centric allocates each user’s contribution $\beta$ across artists in proportion to that user’s streaming shares. It gives artist $i$ the revenue
$R_{i}^{uc}=\sum_{j\in M}\pi_{ij}\beta.$
Let $\overline{\lambda}\equiv\frac{1}{m}\sum_{j\in M}\lambda_{j}$ denote the average number of streams per user, and $d_{i}\equiv\frac{\sum_{j\in M}\lambda_{j}\pi_{ij}}{\sum_{j\in M}\lambda_{j}}$ denote artist $i$’s share of total streams. Then the pro-rata payment can be written as $R_{i}^{pr}=d_{i}m\beta$. Assume $d_{i}<1$ for all $i\in N$ to exclude the trivial case of a monopolist artist.
The model is a non-cooperative game in which the only strategic agents are artists. Each artist $i\in N$ chooses a level of fraud activity $x_{i}\geq 0$, and let $x\in\mathbb{R}_{+}^{N}$ denote the strategy profile. Fraud activity has unit cost $\delta>1$, which covers the subscription fee. Each unit of fraud activity creates a new user who generates $\lambda_{0}$ fake streams on the cheating artist. The parameter $\lambda_{0}$ represents the maximum number of streams per fake user that the platform cannot detect or remove; we refer to $\lambda_{0}$ as the fraud technology, as it captures the platform’s ability to detect fake streams.
Given $x$, the net gain of artist $i$ under pro-rata is
$G_{i}^{pr}(x)=\frac{m\overline{\lambda}d_{i}+\lambda_{0}x_{i}}{m\overline{\lambda}+\lambda_{0}\sum_{j\in N}x_{j}}\Bigg(m+\sum_{j\in N}x_{j}\Bigg)\beta-\delta x_{i},$
where the fraction is artist $i$’s revenue share, $(m+\sum_{j\in N}x_{j})\beta$ is the total revenue allocated to artists, and $\delta x_{i}$ is the fraud cost. In the numerator, $m\overline{\lambda}d_{i}=\sum_{k\in M}\lambda_{k}\pi_{ik}$ denotes real streams on artist $i$, and $\lambda_{0}x_{i}$ denotes fake streams generated by artist $i$. In the denominator, $m\overline{\lambda}=\sum_{k\in M}\lambda_{k}$ is total real streams, and $\lambda_{0}\sum_{j\in N}x_{j}$ is total fake streams.
User-centric is immune to fraud (Bergantiños and Moreno-Ternero, 2025; Ghosh et al., 2025). Under user-centric, the net gain of artist $i$ is
$G_{i}^{uc}(x)=\Bigg(\sum_{j\in M}\pi_{ij}+x_{i}\Bigg)\beta-\delta x_{i},$
where a unit of fraud activity increases the cheating artist’s revenue by $\beta$ while costing $\delta$. Thus, fraud is not profitable under user-centric.
To simplify the analysis, let $\xi=\frac{\delta-\beta}{\beta}$ and $V=\frac{\lambda_{0}-\overline{\lambda}}{\xi\overline{\lambda}}$. Define $t_{i}=\frac{\lambda_{0}x_{i}}{m\overline{\lambda}}$ for all $i\in N$, so that, in the simplified model, artists choose the ratio of fake streams to total real streams rather than the level of fraud activity. Define the utility of artist $i\in N$ as $u_{i}(t)=\frac{\lambda_{0}G_{i}^{pr}(x)}{m\overline{\lambda}(\delta-\beta)}$, which yields
$u_{i}(t)=\frac{d_{i}+t_{i}}{1+\sum_{j\in N}t_{j}}V-t_{i}+d_{i}.$ (1)
The last term is constant and therefore does not affect incentives. All information in the model is public.
3 Analysis
3.1 Fraud-free equilibrium under weak fraud technology
In this royalty-sharing game, the most desirable outcome is an equilibrium in which all artists remain honest, thereby avoiding social waste. To address the vulnerability of the pro-rata to fraud, we first show that honesty is a strict dominant strategy when fraud technology is weak.
Lemma 1.
$t_{i}=0$ is a strict dominant strategy of artist $i$ if and only if $\lambda_{0}\leq\overline{\lambda}[1+\xi h(d_{i})]$, where
$h(d_{i})=\left\{\begin{array}[]{ll}\frac{1}{1-d_{i}}&\text{if $d_{i}\in[0,\frac{1}{2}];$}\\ 4d_{i}&\text{if $d_{i}\in(\frac{1}{2},1]$.}\end{array}\right$.
The function $h(\cdot)$ is continuous and strictly increasing. It implies that the decision to engage in fraud depends on the artist’s real streamshare $d_{i}$. An equivalent condition for honesty can be written as
$\frac{\lambda_{0}-\overline{\lambda}}{\overline{\lambda}}\leq\xi h(d_{i}).$
From this expression, $\xi h(\cdot)$ represents a threshold for the deviation of fraud technology from average streaming. Intuitively, the left-hand side measures how closely a fake user resembles a real user: the smaller the deviation, the more subtle the manipulation. If the platform can detect and remove all fake users whose streaming is no greater than that of real users (i.e., LHS $\leq 0$), then fraud is not profitable and all artists remain honest.
We interpret $\xi$ as the fraud premium and decompose it as
$\xi=\frac{1}{\beta}[(\delta-1)+(1-\beta)],$
where $\delta-1$ is the price of the fraud service and $1-\beta$ is the platform’s retained share of the subscription fee. The fraud premium therefore measures the extra cost per fake user relative to the maximum surplus $\beta$ obtainable from fraud through the revenue pool.
As an immediate implication of Lemma 1, the following theorem characterizes the fraud-free equilibrium.
Theorem 1.
$t^{*}=\mathbf{0}$ is the strict dominant strategy of all artists if and only if the fraud technology satisfies
$\lambda_{0}\leq\overline{\lambda}\bigg(1+\frac{\xi}{1-d_{\min}}\bigg),$
where $d_{\min}\equiv\min_{i\in N}d_{i}$.
Theorem 1 shows that fraud can be avoided under sufficiently weak fraud technology. We call the fraud technology weak if $\lambda_{0}\leq\overline{\lambda}(1+\frac{\xi}{1-d_{\min}})$, and strong if $\lambda_{0}>\overline{\lambda}(1+\frac{\xi}{1-d_{\min}})$.
Theorem 1 also suggests two levers for mitigating fraud. First, the platform can adjust the revenue share parameter $\beta$; in the long run, however, such changes may affect participation (Alaei et al., 2022). Second, the platform can strengthen quality checks, e.g., identity verification, to increase the cost of fraud services $\delta$ and hence the premium $\xi$. The threshold is increasing in the pivotal artist’s streamshare $d_{\min}$. However, balancing streamshares among artists can raise the threshold only up to $\overline{\lambda}(1+\frac{n}{n-1}\xi)$, which may be limited in large platforms ($n\to+\infty$).
3.2 Fraud equilibrium under strong fraud technology
When fraud technology is sufficiently strong, the efficient fraud-free equilibrium collapses, and cheating emerges as a rational strategy.
Lemma 2.
If the fraud technology is sufficiently strong, i.e., $\lambda_{0}>\overline{\lambda}(1+\frac{\xi}{1-d_{\min}})$, then there exists a unique pure strategy Nash equilibrium. This equilibrium is a fraud equilibrium.
The fraud equilibrium refers to an equilibrium with positive fraud activity. Once fraud technology exceeds the critical level, honesty is no longer sustainable. In particular, the pivotal artist with the lowest real streamshare $d_{\min}$ has an incentive to cheat.
Theorem 2 characterizes which artists cheat and by how much.
Theorem 2.
In the fraud equilibrium $t^{*}$, there exists a unique non-empty set of dishonest artists $N^{d}\subseteq N$ with cardinality $n_{d}$, such that $t_{i}^{*}>0$ for all $i\in N^{d}$ and $t_{j}^{*}=0$ for all $j\in N\setminus N^{d}$. The dishonest artists are those with low real streamshares.
  1. (a)
    For all dishonest artists $i,j\in N^{d}$ with $t_{i}^{*},t_{j}^{*}>0$, we have $d_{i}+t_{i}^{*}=d_{j}+t_{j}^{*}$.
  2. (b)
    Let $T^{*}=\sum_{i\in N^{d}}t_{i}^{*}$. The ratio of total streams to real streams is
    $1+T^{*}=\sqrt{\bigg(\frac{n_{d}-1}{2n_{d}}V\bigg)^{2}+\frac{1}{n_{d}}\Bigg(1-\sum_{i\in N^{d}}d_{i}\Bigg)V}+\frac{n_{d}-1}{2n_{d}}V.$ (2)
Next consider the upper bound on the ratio of total streams to real streams $1+T^{*}$.
Proposition 1.
The upper bound of the total-to-real streams ratio $1+T^{*}$ is $V=\frac{\lambda_{0}-\overline{\lambda}}{\xi\overline{\lambda}}$.
Recall that $\frac{\lambda_{0}-\overline{\lambda}}{\overline{\lambda}}$ measures the relative advantage obtained from fraud, and $\xi$ is the fraud premium, measuring the percentage extra cost. Thus, $V$ summarizes the relative advantage from socially wasteful manipulation.
This result is reassuring because even under extreme conditions, total fraud cannot diverge without bound. It suggests that axioms based on unlimited strategic manipulation may be overly strong in this setting (Bergantiños and Moreno-Ternero, 2025; Ghosh et al., 2025): even the summation of all artists’ fraud activities is bounded.
Because dishonest artists are those with low real streamshares, the case in which all artists engage in fraud is referred to as the worst fraud equilibrium. In this case, $N^{d}=N$ and $n_{d}=n$. The ratio of total streams to real streams in the worst fraud equilibrium is $\frac{n-1}{n}V$, since $\sum_{i\in N}d_{i}=1$. In a large platform ($n\to+\infty$), the ratio of fake streams to real streams approaches the maximum V1V-1.
The following corollary provides an equivalent condition for when the worst case occurs.
Corollary 1.
The fraud equilibrium is the worst if and only if
$d_{\mathrm{max}}<\frac{n-1}{n^{2}}V,$
where $d_{\mathrm{max}}\equiv\max_{i\in N}d_{i}$.
This corollary links the largest real streamshare to the worst-case condition: the artist with the largest real streamshare is the least likely to cheat. If $d_{\mathrm{max}}$ is bounded away from zero and entry is free in a large platform, the worst case will not occur.
Now consider a partial fraud equilibrium, where not all artists engage in fraud. By Theorem 2, cheating occurs among low-streamshare artists, and equilibrium fraud is decreasing in real streamshare. Hence there exists a cutoff $d^{*}$ such that artists with $d_{i}<d^{*}$ belong to $N^{d}$, while those with $d_{i}\geq d^{*}$ remain honest.
Corollary 2.
There exists a cutoff $d^{*}$ distinguishing honest artists $j\in N\setminus N^{d}$ with $d_{j}\geq d^{*}$ from dishonest artists $i\in N^{d}$ with $d_{i}<d^{*}$ in a partial fraud equilibrium. That is,
$d^{*}=\frac{1}{n_{d}}\Bigg(\sum_{i\in N^{d}}d_{i}+T^{*}\Bigg).$ (3)
In a large platform with free entry, the cutoff $d^{*}$ approaches zero because the ratio of total fake streams to real streams is bounded. Thus, only artists with negligible real streams, who would otherwise earn almost nothing, choose to engage in fraud. At the same time, the ratio of fake streams to real streams may still approach the upper bound V1V-1.
To summarize, when fraud technology is strong, cheating is confined to low-streamshare artists and aggregate fake streams are bounded. In a large platform with free entry, the worst fraud equilibrium does not occur; instead, only artists with negligible royalties engage in fraud.
3.3 Fairness comparison
So far we have studied equilibrium outcomes under weak and strong fraud technology. The fraud-free equilibrium is efficient but requires sufficiently weak fraud technology, while the fraud equilibrium features bounded social waste. The efficiency comparison is therefore clear: strong fraud technology leads to an inefficient outcome. However, efficiency is not the only criterion.
From the perspective of fairness, we show that fraud may sometimes improve the distribution of payoffs. The fairness criterion considered here is egalitarianism: social welfare depends on the agent with the lowest utility.
Definition 1.
Fix $N$, and let $u$ and $v$ denote two utility profiles. We say that $u$ is fairer than $v$ if $\min_{i\in N}u_{i}>\min_{i\in N}v_{i}$.
We compare the utility profile in the fraud equilibrium with that in the fraud-free situation, where utilities depend only on artists’ real streamshares.
Proposition 2.
In a fraud equilibrium, artists’ utilities have the same ranking as in the fraud-free situation and therefore coincide with the ranking of real streamshares. The utility profile in a fraud equilibrium is fairer than that in the fraud-free situation if and only if
$1+T^{*}<V-\sqrt{d_{\min}V(V-1)}.$
In the fraud equilibrium, the utilities of honest artists decrease because their revenues depend only on real streamshare, while total streams increase due to fraud. Fraud redistributes payoffs from high-share artists to low-share artists by diluting stream value. This redistribution can raise the minimum utility and thereby improve egalitarian fairness. Combining Proposition 2 with Proposition 1 yields an equivalent condition for when the fraud equilibrium is fairer.
Corollary 3.
The worst fraud equilibrium is always fairer than the fraud-free situation if and only if
$d_{\min}<\frac{1}{n^{2}}\frac{V}{V-1}.$
In particular, with free entry ($d_{\min}=0$), the fraud equilibrium is always fairer than the fraud-free situation under this criterion.
4 Discussion on alternative methods
4.1 Parametric weighted rule
From the analysis above, we acknowledge that efficiency would be lost under a strong fraud technology. To address this issue and restore a fraud-free equilibrium through alternative methods, it is not necessary to abandon pro-rata entirely in favor of user-centric as suggested by Dimont (2018). Instead, we propose a parametric weighted rule to share the revenue.
The rule is a convex combination of pro-rata and user-centric, designed to balance the platform’s preference for pro-rata with its limited ability to detect fraud. The revenue paid to artist $i$ under the weighted rule with parameter $\alpha\in(0,1]$ is
$R_{i}^{pw}(\alpha)=\alpha R_{i}^{pr}+(1-\alpha)R_{i}^{uc},$
where a higher $\alpha$ moves the rule closer to pro-rata. The corresponding net gain is
$G_{i}^{pw}(x;\alpha)=\frac{m\overline{\lambda}d_{i}+\lambda_{0}x_{i}}{m\overline{\lambda}+\lambda_{0}\sum_{j\in N}x_{j}}\Bigg(m+\sum_{j\in N}x_{j}\Bigg)\beta\alpha+\Bigg(\sum_{k\in M}\pi_{ik}+x_{i}\Bigg)\beta(1-\alpha)-\delta x_{i}.$
Consider the same game with strategy vector $t$, and define the utility of artist $i\in N$ as $u_{i}^{pw}(t)=\frac{\lambda_{0}G_{i}^{pw}(x)}{m\overline{\lambda}(\delta-\beta)}$. Then
$u_{i}^{pw}(t;\alpha)=\frac{d_{i}+t_{i}}{1+\sum_{j\in N}t_{j}}V\alpha-t_{i}+d_{i}\alpha+\frac{\lambda_{0}}{n\overline{\lambda}}\sum_{k\in M}\pi_{ik}(1-\alpha),$ (4)
where the last two terms do not depend on artist $i$’s choice of $t_{i}$.
Repeating the earlier analysis under this rule yields a generalization of Theorem 1.
Theorem 3.
Under the weighted rule with parameter $\alpha$, $t^{*}=\mathbf{0}$ is the strict dominant strategy of all artists if and only if $\lambda_{0}\leq\overline{\lambda}[1+\frac{\xi}{\alpha(1-d_{\min})}]$.
Theorem 3 shows that the threshold for fraud-proofness increases as $\alpha$ decreases. When $\alpha=1$, the condition reduces to the original pro-rata threshold. As $\alpha\to 0$, the threshold approaches infinity, reflecting the fact that user-centric is inherently immune to fraud.
Theorem 3 also allows us to address the dual question: given a strong fraud technology $\lambda_{0}>\overline{\lambda}(1+\frac{\xi}{1-d_{\min}})$, what values of $\alpha$ guarantee efficiency? The following proposition provides an answer.
Proposition 3.
To guarantee a fraud-free equilibrium under a strong fraud technology, the parameter $\alpha$ must satisfy
$\alpha\leq\frac{1}{(1-d_{\min})V}.$
This condition provides a practical guideline: platforms can mitigate fraud risk by reducing $\alpha$, thereby shifting the rule toward user-centric. The required adjustment depends on $V$, which relates to $\lambda_{0}$, $\overline{\lambda}$, $\delta$, and $\beta$, and on the lowest streamshare $d_{\min}$. In this sense, the proposition identifies a family of weighted rules that preserve efficiency under technological constraints.
4.2 Spotify’s modernized royalty system
At the end of 2023, Spotify announced several new policies to “better support those most dependent on streaming revenues as part of their livelihood.” [^1] The policies include charging for artificial streaming, tracking monetization eligibility, and regulating noise recordings. For more details, see https://artists.spotify.com/en/blog/modernizing-our-royalty-system The qualification policy is our focus: it sets a stream threshold that a track must meet to be eligible to receive royalties. In particular, only eligible tracks, those reached at least 1,000 streams in the past year, enter the pro-rata denominator.
The policy nonetheless affects fraud incentives. Adapting the policy to our model, let $\widehat{\lambda}$ denote the qualification threshold of streams, and let $\widehat{d}=\frac{\widehat{\lambda}}{m\overline{\lambda}}$ denote the corresponding streamshare threshold. We treat each track as an artist for simplicity. An artist $i\in N$ shares revenue if $d_{i}+t_{i}\geq\widehat{d}$; otherwise, the artist pays the fraud cost and receives zero from the revenue pool. Hence any strategy $t_{i}\in(0,\widehat{d}-d_{i})$ is strictly dominated by 0 for artists with $d_{i}<\widehat{d}$, since it incurs cost without qualifying for payment.
We argue that this policy may strictly increase fake streams and hurt all artists. To illustrate this possibility, we focus on a reasonable range of thresholds $\widehat{d}$ that are not too high: an overly demanding threshold could affect artists who rely on streaming revenue, and a full characterization for arbitrary thresholds is beyond the scope of this paper. Within the range we consider, dishonest artists in the baseline equilibrium still cheat.
Consider a value of $V$ for which the baseline model yields a fraud equilibrium $t^{*}$ with a set of dishonest artists $N^{d}$ and a cutoff $d^{*}$. If the platform chooses a low qualification threshold $\widehat{d}\leq d^{*}$, the policy fails: dishonest artist $i\in N^{d}$ with $d_{i}<\widehat{d}\leq d^{*}$ still choose $t_{i}^{*}=d^{*}-d_{i}\geq\widehat{d}-d_{i}$, which is unaffected by eliminating dominated strategies.
Suppose instead that the platform chooses a threshold $\widehat{d}$ slightly above $d^{*}$. [^2] The range of slightly-high $\widehat{d}$ is non-empty, as shown in Appendix A.8. In this range, all artists cheat at least the same as in the baseline equilibrium, and no artist is eliminated. Then a low-streamshare artist chooses either $t_{i}=0$ or $t_{i}\geq\widehat{d}-d_{i}$. In a new equilibrium where all initial cheaters still cheat, their best response is the just-qualify strategy $\widehat{d}-d_{i}$, because the resulting increase in total fake streamshare reduces the marginal return to additional fraud. Let $\widehat{t}$ be defined by $\widehat{t}_{i}=\max\{\widehat{d}-d_{i},0\}$ for all $i\in N$, and $\widehat{T}$ denote aggregate fraud under $\widehat{t}$. The strategy profile $\widehat{t}$ is the equilibrium under threshold $\widehat{d}$. In this equilibrium, artist $i$ with $\widehat{d}>d_{i}$ will cheat more, i.e., $\widehat{t}_{i}>t_{i}^{*}$; and artist $i$ with $\widehat{d}\leq d_{i}$ will keep honest, i.e., $\widehat{t}_{i}=t^{*}$, and thus $\widehat{T}-\widehat{t}_{i}>T^{*}-t_{i}^{*}$. Consequently, for all $i\in N$,
[…]
where two inequalities cannot both hold with equality, so at least one is strict. Thus all artists are worse off. In addition, it reduces both efficiency and fairness.
5 Concluding remarks
This paper develops a tractable model of strategic manipulation by artists on music streaming platforms under the pro-rata revenue sharing rule. The framework is flexible enough to accommodate evolving detection methods and increasingly sophisticated fraud strategies, including those enabled by machine learning and advanced identification technologies. It shows that pro-rata can be robust to fraud when fraud technology is weak, yielding an efficient fraud-free equilibrium. When fraud technology is strong, a unique fraud equilibrium exists, but aggregate fake streams remain bounded because increased total streams dilute the value of manipulation. Fraud may also compress payoffs and improve egalitarian fairness despite efficiency losses.
Appendix A Proofs
A.1 Proof of Lemma 1
Let $\tau_{i}=\sum_{j\in N\setminus\{i\}}t_{j}$ denote the summation of all agents’ strategy except agent $i$. Rewrite the utility function as $v_{i}(t_{i},\tau_{i})$. That is,
$v_{i}(t_{i},\tau_{i})=\frac{d_{i}+t_{i}}{1+t_{i}+\tau_{i}}V-t_{i}+d_{i}.$ (5)
To show $t_{i}=0$ is a strict dominant strategy of agent $i$, one needs to show that for all $\tau_{i}\geq 0$, it holds for all $t_{i}>0$ that $v_{i}(0,\tau_{i})>v_{i}(t_{i},\tau_{i})$. It is
$\bigg(\frac{d_{i}}{1+\tau_{i}}-\frac{d_{i}+t_{i}}{1+t_{i}+\tau_{i}}\bigg)V+t_{i}>0,$ (6)
which is equivalent to
$(1+\tau_{i})t_{i}+(1+\tau_{i})^{2}+(d_{i}-1-\tau_{i})V>0.$ (7)
So the statement is equivalent to that for all $\tau_{i}\geq 0$,
$(1+\tau_{i})^{2}+(d_{i}-1-\tau_{i})V\geq 0,$
which is equivalent to
$\bigg(\tau_{i}+1-\frac{1}{2}V\bigg)^{2}\geq\frac{1}{4}V^{2}-d_{i}V.$ (8)
Case 1: $V\leq 2$. (8) holds for all $\tau_{i}\geq 0$ would be equivalent to
$\bigg(1-\frac{1}{2}V\bigg)^{2}\geq\frac{1}{4}V^{2}-d_{i}V.$
It is
$V\leq\frac{1}{1-d_{i}}.$
If $d_{i}\leq\frac{1}{2}$, there is $\frac{1}{1-d_{i}}\leq 2$. So $V\leq\frac{1}{1-d_{i}}$. If $d_{i}>\frac{1}{2}$, there is $\frac{1}{1-d_{i}}>2$. So $V\leq 2$.
Case 2: $V>2$. (8) holds for all $\tau_{i}\geq 0$ would be equivalent to
$0\geq\frac{1}{4}V^{2}-d_{i}V.$
That is,
$V\leq 4d_{i}.$
If $d_{i}\leq\frac{1}{2}$, there is $V\leq 4d_{i}\leq 2<V$, which is not the case. If $d_{i}>\frac{1}{2}$, there is $2<V\leq 4d_{i}$. So $V\leq 4d_{i}$.
To sum up, if $d_{i}\leq\frac{1}{2}$, $V\leq\frac{1}{1-d_{i}}$; if $d_{i}>\frac{1}{2}$, $V\leq 4d_{i}$.
A.2 Proof of Lemma 2
To show Lemma 2, three steps are constructed. First step shows the existence of Nash equilibrium for $V>0$. Second step shows the uniqueness of Nash equilibrium for $V\geq 1$. Third step shows that $t^{*}=\mathrm{0}$ is not a Nash equilibrium when $V>\frac{1}{1-d_{\min}}$.
Step 1 (Existence): Assume $V>0$. The derivative of agent $i$’s utility function (5) is
$\frac{\partial v_{i}(t_{i},\tau_{j})}{\partial t_{i}}=\frac{1-d_{i}+\tau_{i}}{(1+t_{i}+\tau_{i})^{2}}V-1,$ (9)
which approaches 1-1 as $t_{i}\rightarrow+\infty$. The second-order derivative is
$\frac{\partial^{2}v_{i}(t_{i},\tau_{j})}{\partial t_{i}^{2}}=-2\frac{(1-d_{i}+\tau_{i})}{(1+t_{i}+\tau_{i})^{3}}V,$
which is negative. So the utility function is strictly concave.
Now prove the following claim: There exists an upper bound $\overline{t}_{i}\geq 0$ such that for all $\tau_{i}\geq 0$, $\frac{\partial v_{i}(t_{i},\tau_{i})}{\partial t_{i}}<0$ holds for all $t_{i}>\overline{t}_{i}$. Let $\overline{t}_{i}=\frac{1}{2}V$, then it is to show that for all $\tau_{i}\geq 0$, all $t_{i}>\frac{V}{2}$,
$(1-d_{i}+\tau_{i})V<(1+t_{i}+\tau_{i})^{2}.$
It is equivalent to that for all $\tau_{i}\geq 0$,
$(1-d_{i}+\tau_{i})V\leq\bigg(1+\frac{1}{2}V+\tau_{i}\bigg)^{2}.$
That is,
$(1+\tau_{i})^{2}+\frac{1}{4}V^{2}+d_{i}V\geq 0,$
which holds for all $\tau_{i}\geq 0$.
So all strategies exceeding $\frac{1}{2}V$ are strictly dominated. Therefore the strategy set of agent $i$ can be restricted into $0,\frac{1}{2}V]$, which is compact and convex. Combining with the continuity and convexity of utility function, Nash theorem ensures the existence of Nash equilibrium.
Step 2 (Uniqueness): Take $V\geq 1$. Let $t^{*}$ denote the Nash equilibrium and $T^{*}=\sum_{i\in N}t^{*}_{i}$ denote the aggregate strategy in equilibrium. Existence of Nash equilibrium implies that there exists a set of agents $N^{d}\subseteq N$ with positive strategy in equilibrium. Due to concavity of utility function, for all $i\in N^{d}$, there are $t^{*}_{i}>0$ and $\frac{\partial v_{i}(t^{*}_{i},\tau^{*}_{i})}{\partial t_{i}}=0$, which, from (9), is equivalent to
$d_{i}+t^{*}_{i}=1+T^{*}-\frac{1}{V}(1+T^{*})^{2}.$ (10)
Also for all $j\in N\setminus N^{d}$, there are $t^{*}_{j}=0$ and
$d_{j}\geq 1+T^{*}-\frac{1}{V}(1+T^{*})^{2}.$ (11)
So the positivity of $t^{*}_{i}$ is determined by $1+T^{*}-\frac{1}{V}(1+T^{*})^{2}$ and $d_{i}$.
Consider a helper function $f_{i}(T^{*})=\frac{d_{i}+t^{*}_{i}}{1+T^{*}}$ for all $i\in N$, of which the form depends on $d_{i}$.
Case 1: $d_{i}\geq\frac{1}{4}V$. The following inequality always holds:
$\bigg(1+T^{*}-\frac{1}{2}V\bigg)^{2}\geq\frac{1}{4}V^{2}-d_{i}V,$
which is equivalent to (11). So in equilibrium agent $i$ would always take strategy $t^{*}_{i}=0$. Thus $f_{i}(T^{*})=\frac{d_{i}}{1+T^{*}}$, which is continuous and strictly decreasing on $\mathbb{R}_{+}$.
Case 2: $d_{i}<\frac{1}{4}V$. From (11), we know that $t^{*}_{i}=0$ if either $1+T^{*}\leq\frac{1}{2}V-\sqrt{\frac{1}{4}V^{2}-d_{i}V}$, or $1+T^{*}\geq\frac{1}{2}V+\sqrt{\frac{1}{4}V^{2}-d_{i}V}$. Otherwise, $d_{i}+t^{*}_{i}=1+T^{*}-\frac{1}{V}(1+T^{*})^{2}$. So that
$f_{i}(T^{*})$ $=\left\{\begin{array}[]{ll}\frac{d_{i}}{1+T^{*}}&\text{if $T^{*}\leq\frac{1}{2}V-\sqrt{\frac{1}{4}V^{2}-d_{i}V}-1$;}\\ 1-\frac{1}{V}(1+T^{*})&\text{if $\frac{1}{2}V-\sqrt{\frac{1}{4}V^{2}-d_{i}V}-1<T^{*}<\frac{1}{2}V+\sqrt{\frac{1}{4}V^{2}-d_{i}V}-1$;}\\ \frac{d_{i}}{1+T^{*}}&\text{if $T^{*}\geq\frac{1}{2}V+\sqrt{\frac{1}{4}V^{2}-d_{i}V}-1$,}\end{array}\right$.
which is continuous and strictly decreasing on $\mathbb{R}_{+}$.
Let $F(T^{*})\equiv\sum_{i\in N}f_{i}(T^{*})$. Since $f_{i}(T^{*})$ is continuous and strictly decreasing on $\mathbb{R}_{+}$ for all $i\in N$, $F(T^{*})$ is continuous and strictly decreasing on $\mathbb{R}_{+}$. Notice that
$1=\frac{1+T^{*}}{1+T^{*}}=\sum_{i\in N}\frac{d_{i}+t^{*}_{i}}{1+T^{*}}=F(T^{*}).$
From $f_{i}(T^{*})=\frac{d_{i}+t^{*}_{i}}{1+T^{*}}\geq\frac{d_{i}}{1+T^{*}}$, there is $F(T^{*})\geq\sum_{i\in N}\frac{d_{i}}{1+T^{*}}=\frac{1}{1+T^{*}}$. So $F(0)\geq 1$. Combining with $F(V-1)=\frac{1}{V}\leq 1$, as well as the constant LHS, it yields a unique aggregate strategy $T^{*}\in[0,V-1]$. Therefore the uniqueness can be derived by unique $T^{*}$ as well as constant $d_{i}$ for all $i\in N$.
Step 3 (Non-existence at zero): This step shows that $t^{*}=\mathbf{0}$ is not a Nash equilibrium when $V>\frac{1}{1-d_{\min}}$ by contradiction. Suppose $t^{*}=\mathbf{0}$ is a Nash equilibrium. Consider the pivotal agent $i$ with $d_{i}=d_{\min}$. There is $v_{i}(0,0)\geq v_{i}(t_{i},0)$ for all $t_{i}>0$. Substituting $\tau_{i}=0$ into (7), it is,
$t_{i}+1\geq(1-d_{\min})V.$ (12)
One can find a strategy
$\widetilde{t}_{i}=\sqrt{(1-d_{\min})V}-1>0,$
as $V>\frac{1}{1-d_{\min}}\geq 1$. Substituting $\widetilde{t}_{i}$ into (12), there is
$\sqrt{(1-d_{\min})V}\geq(1-d_{\min})V,$
which is equivalent to $(1-d_{\min})V\leq 1$ contradicting $V>\frac{1}{1-d_{\min}}$. Thus $t^{*}=\mathbf{0}$ is not a Nash equilibrium.
To sum up, if $V>\frac{1}{1-d_{\min}}$, there exists a unique fraud equilibrium, where at least one agent takes positive strategy.
A.3 Proof of Theorem 2
From Lemma 2, there exists a unique fraud equilibrium. Thus there is a unique non-empty subset of agents $N^{d}\subseteq N$ taking positive strategies in equilibrium. Following lemma shows that the agents can be separated according to real streamshare $d_{i}$ and those who cheat are those with less real streamshare.
Lemma 3.
Suppose $t^{*}$ is a fraud equilibrium, and $N^{d}$ is a set of agents who take positive strategies in equilibrium. Then for all pair $i,j\in N$:
  1. (a)
    If $i\in N^{d}$ and $d_{j}\leq d_{i}$, then $j\in N^{d}$;
  2. (b)
    If $i\in N\setminus N^{d}$ and $d_{j}\geq d_{i}$, then $j\in N\setminus N^{d}$.
Proof. (a) Given $i\in N^{d}$ and $d_{j}\leq d_{i}$, suppose $j\in N\setminus N^{d}$. There is $t^{*}_{i}>0$, $t^{*}_{j}=0$, $\frac{\partial v_{i}(t_{i}^{*},\tau_{i}^{*})}{\partial t_{i}}=0$, and $\frac{\partial v_{j}(t_{j}^{*},\tau_{j}^{*})}{\partial t_{j}}\leq 0$. From (10), it implies
$1+T^{*}-\frac{1}{V}(1+T^{*})^{2}\leq d_{j}+t^{*}_{j}=d_{j}\leq d_{i}<d_{i}+t^{*}_{i}=1+T^{*}-\frac{1}{V}(1+T^{*})^{2},$
which is a contradiction. Thus $j\in N^{d}$.
(b) Conversely, given $i\in N\setminus N^{d}$ and $d_{j}\geq d_{i}$, suppose $j\in N^{d}$. There is $t^{*}_{i}=0$, $t^{*}_{j}>0$, $\frac{\partial v_{i}(t_{i}^{*},\tau_{i}^{*})}{\partial t_{i}}\leq 0$ and $\frac{\partial v_{j}(t_{j}^{*},\tau_{j}^{*})}{\partial t_{j}}=0$. From (10), it implies
$1+T^{*}-\frac{1}{V}(1+T^{*})^{2}=d_{j}+t^{*}_{j}>d_{j}\geq d_{i}=d_{i}+t^{*}_{i}\geq 1+T^{*}-\frac{1}{V}(1+T^{*})^{2},$
which is a contradiction. Thus $j\in N\setminus N^{d}$.
Next study the fraud equilibrium with set $N^{d}$. For all $i,j\in N^{d}$, repeating (10) gives
$d_{i}+t_{i}^{*}=1+T^{*}-\frac{1}{V}(1+T^{*})^{2}=d_{j}+t^{*}_{j}.$
Notice that $T^{*}=\sum_{i\in N^{d}}t^{*}_{i}$. Summing up (10) for all $i\in N^{d}$, it is
$\sum_{i\in N^{d}}d_{i}+T^{*}=n_{d}\bigg[(1+T^{*})-\frac{1}{V}(1+T^{*})^{2}\bigg].$ (13)
It is equivalent to
$\bigg(1+T^{*}-\frac{n_{d}-1}{2n_{d}}V\bigg)^{2}=\bigg(\frac{n_{d}-1}{2n_{d}}V\bigg)^{2}+\frac{1}{n_{d}}\Bigg(1-\sum_{i\in N^{d}}d_{i}\Bigg)V,$
where the right-hand side (RHS) is non-negative. Thus
$1+T^{*}=\sqrt{\bigg(\frac{n_{d}-1}{2n_{d}}V\bigg)^{2}+\frac{1}{n_{d}}\Bigg(1-\sum_{i\in N^{d}}d_{i}\Bigg)V}+\frac{n_{d}-1}{2n_{d}}V.$
A.4 Proof of Proposition 1
Let $r(N^{d})=1+T^{*}$ a function of set of agents $N^{d}$ with positive strategies in equilibrium. That is,
$r(N^{d})=\sqrt{\bigg(\frac{n_{d}-1}{2n_{d}}V\bigg)^{2}+\frac{1}{n_{d}}\Bigg(1-\sum_{i\in N^{d}}d_{i}\Bigg)V}+\frac{n_{d}-1}{2n_{d}}V.$
Consider two helper functions:
$\underline{r}(s)$ $=\frac{s-1}{s}V;$
$\overline{r}(s)$ $=\sqrt{\bigg(\frac{s-1}{2s}V\bigg)^{2}+\frac{1}{s}V}+\frac{s-1}{2s}V,$
where $s$ is treated as a continuous variable. Notice that for all possible $N^{d}$ in fraud equilibrium, there is
$\underline{r}(n_{d})\leq r(N^{d})\leq\overline{r}(n_{d}).$ (14)
It’s obvious that $\underline{r}(s)$ is strictly increasing in $s$. Notice that $\overline{r}$ is a positive root of the following equation:
$\overline{r}^{2}-\frac{s-1}{s}V\overline{r}-\frac{1}{s}V=0,$
while another root is negative. Let $F(\overline{r},s)=\overline{r}^{2}-\frac{s-1}{s}V\overline{r}-\frac{1}{s}V$. Notice that F(1,s)=1V<0F(1,s)=1-V<0, and $F(\overline{r},s)\rightarrow+\infty\ (\overline{r}\rightarrow+\infty)$. So that $\overline{r}>1$. The derivative of $\overline{r}(s)$ is
$\overline{r}^{\prime}(s)$ $=\frac{1}{2s^{2}}V\Bigg[\frac{\frac{s-1}{2s}V-1}{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}}+1\Bigg].$
$=\frac{1}{2s^{2}}V\frac{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}+\frac{s-1}{2s}V-1}{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}}$
$=\frac{1}{2s^{2}}V\frac{\overline{r}(s)-1}{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}},$
which is positive. So $\overline{r}(s)$ is strictly increasing. Also,
$\overline{r}(n)=\sqrt{\bigg(\frac{n-1}{2n}V\bigg)^{2}+\frac{1}{n}V}+\frac{n-1}{2n}V,$
which approaches its upper bound $V$ as $n\rightarrow+\infty$. Combining with the fact that $\underline{r}(n)$ also approaches $V$ as well. The upper bound of $1+T^{*}$ is thus $V$.
A.5 Proofs of Corollaries 1, 2, and 3
Proof of Corollary 1. From statement (a) of Theorem 2, equilibrium strategy of agent $i\in N^{d}$ is decreasing in $d_{i}$. All agents taking positive strategies is thus equivalent to that the agent with $d_{\mathrm{max}}$ takes positive strategy. Suppose agent $i$ is the agent with $d_{i}=d_{\mathrm{max}}$. From (10), the equilibrium strategy of agent $j$ is
$d_{\mathrm{max}}+t_{j}^{*}=1+T^{*}-\frac{1}{V}(1+T^{*})^{2}.$
Then $t_{j}^{*}>0$ is equivalent to
$d_{\mathrm{max}}<1+T^{*}-\frac{1}{V}(1+T^{*})^{2}.$
Notice that $1+T^{*}=\frac{n-1}{n}V$ holds in the worst fraud equilibrium. So it is equivalent to
$d_{\mathrm{max}}<\frac{n-1}{n}V-\frac{1}{V}(\frac{n-1}{n}V)^{2}=\frac{n-1}{n^{2}}V.$
Thus the equivalent condition is obtained.
Proof of Corollary 2. One can find a threshold
$d^{*}$ $=1+T^{*}-\frac{1}{V}(1+T^{*})^{2}.$ (15)
Now check two cases. For all $i\in N^{d}$ with $t_{i}^{*}>0$. From (10), there is
$1+T^{*}-\frac{1}{V}(1+T^{*})^{2}=d_{i}+t^{*}_{i}>d_{i},$
which is equivalent to $d_{i}<d^{*}$. Also, for all $j\in N\setminus N^{d}$ with $t^{*}_{j}=0$. From $\frac{\partial v_{j}(t_{j}^{*},\tau_{j}^{*})}{\partial t_{j}}\leq 0$, there is
$1+T^{*}-\frac{1}{V}(1+T^{*})^{2}\leq d_{j}+t^{*}_{j}=d_{j},$
which is exactly $d_{j}\geq d^{*}$. Substituting (13) into (15), formula shown in corollary can be derived.
Proof of Corollary 3. From Theorem 2, as well as $\sum_{i\in N}d_{i}=1$, there is $1+T^{*}=\frac{n-1}{n}V$. Combining with Proposition 2, the statement holds if and only if
$\frac{n-1}{n}V<V-\sqrt{d_{\min}V(V-1)},$
which is equivalent to the formula.
A.6 Proof of Proposition 2
Consider two utility profiles $u^{*}$ and $u^{0}$, corresponding to that in fraud equilibrium and the fraud-free situation. Obviously, $u_{i}^{0}=u_{i}(\mathbf{0})=d_{i}(1+V)$ for all $i\in N$. From (10) and (15), for all agent $i\in N^{d}$ with positive equilibrium strategy, there is
$d_{j}+t_{j}^{*}=d^{*},$
while for all agent $j\in N\setminus N^{d}$ with $t_{j}^{*}=0$, there is $d_{j}\geq d^{*}$. So that for all agent $i\in N^{d}$,
$u_{i}^{*}=\frac{d^{*}}{1+T^{*}}V-d^{*}+2d_{i}.$ (16)
The first two terms are independent of $i\in N^{d}$. Thus for all agents in $N^{d}$, utilities among profile $u^{*}$ follow the same order as those in $u^{0}$.
Next we show that, in partial fraud equilibrium, the utility of any agent in $N\setminus N^{d}$ would be higher than that of all agents in $N^{d}$. Consider any pair $i\in N^{d}$ and $j\in N\setminus N^{d}$ in partial fraud equilibrium, where $N\setminus N^{d}$ is not empty. Notice that
$u_{j}^{*}=d_{j}\bigg(1+\frac{V}{1+T^{*}}\bigg)=\frac{d_{j}}{1+T^{*}}V-d^{*}+d^{*}+d_{j}>\frac{d^{*}}{1+T^{*}}V-d^{*}+2d_{i}=u_{i}^{*}.$
Thus we can conclude that utilities in the fraud equilibrium follow the same order as those in the fraud-free situation.
To show the fairness property, the comparison between two profiles for the utility of pivotal agent $i$ with $d_{i}=d_{\min}$ is needed. Combining (16) and $u_{i}^{0}=d_{\min}(1+V)$, there is
$u_{j}^{*}-u_{j}^{0}$ $=\frac{d^{*}}{1+T^{*}}V-d^{*}+2d_{\min}-d_{\min}(1+V)$
$=\frac{1}{V}[(1+T^{*})-V]^{2}-d_{\min}(V-1).$
So that the fraud equilibrium is fairer, i.e., $u_{i}^{*}>u_{i}^{0}$, is equivalent to
$[(1+T^{*})-V]^{2}>d_{\min}V(V-1).$
That is, either $1+T^{*}>V+\sqrt{d_{\min}V(V-1)}$, or $1+T^{*}<V-\sqrt{d_{\min}V(V-1)}$. From Proposition 1, there is $1+T^{*}\leq V$. Thus only the later range is kept.
A.7 Proof of Theorem 3
Similar to the proof of Lemma 1. Rewrite the parametric utility function of agent $i$ as follows.
$v_{i}(t_{i},\tau_{i};\alpha)=\frac{d_{i}+t_{i}}{1+t_{i}+\tau_{i}}V\alpha-t_{i}+d_{i}\alpha+\frac{\lambda_{0}}{n\overline{\lambda}}\sum_{k\in M}\pi_{ik}(1-\alpha)$ (17)
$t_{i}=0$ is a strict dominant strategy of agent $i$ is equivalent to that for all $\tau_{i}\geq 0$, $v_{i}(0,\tau_{i};\alpha)>v_{i}(t_{i},\tau_{i};\alpha)$ holds for all $t_{i}>0$. That is,
$\bigg(\frac{d_{i}}{1+\tau_{i}}-\frac{d_{i}+t_{i}}{1+t_{i}+\tau_{i}}\bigg)V\alpha+t_{i}>0,$
which is exactly (6) if $V$ is replaced with $V\alpha$. Repeating the analysis of Lemma 1 gives the equivalent formula. That is, $V\alpha\leq\frac{1}{1-d_{i}}$ if $d_{i}\leq\frac{1}{2}$, and $V\alpha\leq 4d_{i}$ if $d_{i}>\frac{1}{2}$. Thus the theorem can be immediately derived.
A.8 Existence of equilibrium under slightly high threshold
Now verify that there exists such policy $\widehat{d}$ leading to a equilibrium that all artists are not eliminated. The focus is on the pivotal artist $i$, who is most likely to switch from just-qualifying to honesty as $\widehat{d}$ increases. The pivotal artist will stay and play the just-qualify strategy when $u_{i}(\widehat{t})\geq 0$. Notice that $1+T^{*}<1+\widehat{T}\leq 1+T^{*}+n(\widehat{d}-d^{*})$. There is
$u_{i}(\widehat{t})=\frac{\widehat{d}}{1+\widehat{T}}V-\widehat{d}+2d_{\min}\geq\frac{d^{*}+(\widehat{d}-d^{*})}{1+T^{*}+n(\widehat{d}-d^{*})}V-d^{*}-(\widehat{d}-d^{*})+2d_{\min}.$
Let $\theta\equiv\widehat{d}-d^{*}>0$ and consider the following helper function of the RHS.
$H(\theta)=\frac{d^{*}+\theta}{1+T^{*}+n\theta}V-d^{*}-\theta+2d_{\min}.$
Its derivative is
$H^{\prime}(\theta)=\frac{1+T^{*}-nd^{*}}{(1+T^{*}+n\theta)^{2}}V-1.$
If $\frac{1+T^{*}}{nd^{*}}<\frac{V}{V-1}$, then $H^{\prime}(\theta)<0$ always hold. Let $\theta^{*}=0$ in this case. If $\frac{1+T^{*}}{nd^{*}}\geq\frac{V}{V-1}$, let $\theta^{*}=\max\{\frac{1}{n}[\sqrt{(1+T^{*}-nd^{*})V}-(1+T^{*})],0\}$, then $H^{\prime}(\theta)<0$ on $(\theta^{*},+\infty)$, and $H^{\prime}(\theta)>0$ on $(0,\theta^{*})$ if $\theta^{*}>0$.
From two cases, we have $H(\theta^{*})\geq H(0)=u_{i}(t^{*})>u_{i}(0,t^{*}_{-i})\geq 0$, and $H(V+d_{\min})=(\frac{d^{*}+V+d_{\min}}{1+T^{*}+nV+nd_{\min}}-1)V-(d^{*}-d_{\min})<0$. Thus there exists a $\theta^{**}\in(\theta^{*},V+d_{\min})$ such that $H(\theta^{**})=0$. Combining that $H(\theta)$ is strictly increasing on $(0,\theta^{*})$ for positive $\theta^{*}$, and strictly decreasing on $(\theta^{*},V+d_{\min})$, it has $H(\theta)\geq 0$ for all $\theta\in[0,\theta^{**}]$. Thus we can find a non-empty range for $\widehat{d}\in(d^{*},d^{*}+\theta^{**}]$, such that $u_{i}(\widehat{t})\geq H(\theta^{**})\geq 0$, verifying the existence of new equilibrium.